FROM A SQUARE PIECE OF CARDBOARD WITH WIDTH x INCHES, A SQUARE OF WIDTH x-3 INCHES IS REMOVED FROM THE CENTER. WHAT IS THE AREA OF THE REMAINING PIECE AS A FUNCTION OF x.
Answer by uma(370)

Width of the square piece of cardboard = x inches
Area of the square piece = width x width
= x^2 sq inches
Width of the removed square = x-3 inches.
Area of the removed square = (x-3)^2 sq inches
Area of the remaining piece = x^2 - (x-3)^2
==> = x^2 - (x^2 -6x + 9) [expand (x-3)^2]
= x^2 - x^2 + 6x - 9
= 6x - 9
Area of the remaining piece as a function of x = 6x - 9
Good Luck!!!